MEGHÍVÓ

2001. szeptember 10–én, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

Kenneth KUNEN (University of Wisconsin, USA)

BOHR TOPOLOGIES

című előadása.

If G is any (discrete) group, it embeds densely in its Bohr compactification (bG), and this embedding induces a topology on G called the Bohr topology (G#). This subject has its roots in the theory of group representations from the 1930s, but has recently given rise to a number of questions in set-theoretic topology.

For non-abelian groups, there are still many questions about the cardinality of bG and the weight of G# (e.g., does it have a countable base?). For abelian groups, the cardinality of bG and the weight of G# depend only on the cardinality of G, and research has focused on the extent to which the topology of G# determines the algebraic structure of G.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!

2001. szeptember 10–én, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

Kenneth KUNEN (University of Wisconsin, USA)

BOHR TOPOLOGIES

című előadása.

If G is any (discrete) group, it embeds densely in its Bohr compactification (bG), and this embedding induces a topology on G called the Bohr topology (G). This subject has its roots in the theory of group representations from the 1930s, but has recently given rise to a number of questions in set-theoretic topology.

For non-abelian groups, there are still many questions about the cardinality of bG and the weight of G (e.g., does it have a countable base?). For abelian groups, the cardinality of bG and the weight of G depend only on the cardinality of G, and research has focused on the extent to which the topology of G determines the algebraic structure of G.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!